# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ IV-A Evolution of 3GPP Standards for Positioning
1G networks were based on analog communication and lacked standardized positioning functionalities. Positioning was primarily achieved through signal strength measurements, such as RSS, which were used for optimizing BS selection, channel allocation, and handovers. Proprietary solutions, such as TDOA and AOA in advanced mobile phone system (AMPS) signals, supported emergency services and intelligent transportation systems [1, 170]. However, the overall positioning accuracy remained low.
2G networks introduced digital communications, and positioning technologies were initially standardized in the global system for mobile communications (GSM) [1], including cell-ID, time advance, enhanced observed time difference (E-OTD), and assisted GPS (A-GPS) [171]. These technologies significantly improved positioning accuracy, supported emergency services (such as E911), and laid the foundation for the development of subsequent positioning technologies.
Positioning technologies advanced further in 3G networks with the adoption of universal mobile telecommunications system (UMTS) and CDMA2000 standards. UMTS introduced techniques such as OTDOA, uplink time difference of arrival (UTDOA), and A-GPS [172]. radio frequency pattern matching (RFPM) was also employed to optimize performance [170]. CDMA2000 combined advanced forward link trilateration (AFLT) with A-GPS, leveraging network synchronization for efficient positioning [170]. These methods improved positioning accuracy to a range of 25–200 meters, addressing higher performance requirements in various applications.
With the introduction of LTE Release 9, positioning technologies were standardized for the first time. This standardization not only enhanced positioning accuracy but also addressed diverse scenarios and requirements. Positioning became an essential module for ensuring service quality, optimizing networks, and supporting emergency response. Key positioning methods introduced in LTE include [173]:
@@
To support cellular positioning, PRS signals were transmitted in dedicated subframes to minimize interference and enhance OTDOA performance. Furthermore, LTE positioning protocols, such as LTE positioning protocol (LPP), were defined to facilitate efficient communication between mobile terminals and positioning servers.
Subsequently, LTE-advanced (LTE-A) introduced UTDOA, RFPM [174], and enhanced OTDOA methods, improving positioning accuracy through multi-antenna transmission, heterogeneous networks and cross-network measurement [170]. The performance of PRS signals and cell hearability was also optimized. LTE-A also further enhanced positioning capabilities, introducing D2D-assisted positioning, MIMO vertical positioning and Bluetooth/WLAN/barometer positioning to meet positioning needs in multiple scenarios[175].
@@
With the evolution of 5G technologies, positioning capabilities in 3GPP standards have advanced significantly to meet the growing demands of diverse applications. In Release 16, the groundwork for 5G positioning was laid by defining positioning use cases and service requirements [176, 177]. These efforts were aimed at addressing the initial needs for positioning in 5G networks. Subsequent studies focused on enabling new radio (NR) positioning techniques across both FR1 (sub-6 GHz) and FR2 (mmWave) bands. Key aspects included the specification of NR Downlink and Uplink reference signals to support techniques such as Downlink TDOA (DL-TDOA), Downlink AOD (DL-AOD), Uplink TDOA (UL-TDOA), Uplink AOA (UL-AOA), Multi-cell RTT, and E-CID. Additionally, the NR positioning protocol A (NRPPa) was introduced [178], providing a standardized communication framework for positioning in 5G networks.
To meet the demands of emerging 5G applications and vertical industries requiring higher accuracy, lower latency, and improved reliability, 3GPP launched the "Study on NR Positioning Enhancements" in Release 17 [179]. This study focused on solutions for high-accuracy, low-latency positioning that would also improve network and device efficiency. Key advancements included exploring positioning in in-coverage, partial coverage, and out-of-coverage scenarios to support diverse use cases, including sidelink-based positioning for V2X applications [180].
In Release 18, 3GPP further advanced NR positioning with the "Study on expanded and improved NR positioning" [181]. This release introduced multiple enhancements aimed at improving positioning performance and supporting new use cases. To enhance positioning accuracy, it proposed PRS and sounding reference signal (SRS) bandwidth aggregation for intra-band carriers and NR carrier phase measurements. In addition, Release 18 supported low-power high-accuracy positioning and positioning for RedCap UEs. Considering the needs of V2X positioning, sidelink-based positioning was also studied in [181].
@@
For AI/ML aspect, starting with Release 18, 3GPP initiated a study on AI/ML positioning, as outlined in [182]. This study explores the potential of AI/ML to improve positioning accuracy, details the general AI/ML framework, use cases for AI/ML positioning, and evaluation metrics and common KPIs for AI/ML positioning. These developments aim to complement traditional positioning methods by leveraging data-driven models to process large datasets and provide more accurate results.
TABLE I: 5G Positioning performance requirements
@@
Level 1 $10$ m $3$ m N/A N/A $95$ 1 s $\leq 30$ km/h (indoor) $\leq 250$ km/h (rural and urban) $\leq 500$ km/h (trains)
Level 2 $3$ m $99$
Level 3 $1$ m $2$ m
@@
Level 4 $1$ m $2$ m N/A N/A $99.9$ $15$ms $\leq 30$ km/h indoor
Level 6 $0.3$ m $2$ m $10$ms $\leq 30$ km/h indoor
@@ IV-B KPIs for Wireless Positioning
Positioning performance is assessed through a set of critical metrics that ensure the accuracy and reliability of location services. The 5G system is designed to deliver positioning services that meet the performance requirements outlined in Table I. These requirements are inclusive of all UE types, including specialized UEs such as V2X and machine-type communication (MTC) devices.
In 3GPP [177], 5G positioning use cases are divided into 5G positioning service area and 5G enhanced positioning service area. The former includes indoor and outdoor (rural and urban) scenarios. Indoor includes location inside buildings such as offices, hospital, industrial buildings. Outdoor scenarios support the positioning of vehicles and trains at speeds up to 250 km/h and 500 km/h. And the 5G enhanced positioning service area further supports dense urban areas (up to 60 km/h), vehicles in tunnels, and railway positioning.
@@
In addition to horizontal and vertical positioning accuracy, 3GPP Release 19 also introduces positioning service availability and positioning service latency to support mission-critical services. A total of 7 positioning service levels are defined, including 6 levels of absolute positioning accuracy requirements and 1 relative positioning accuracy requirement. In the absolute positioning accuracy requirements, the horizontal accuracy ranges from 10 m to 0.3 m, and the vertical accuracy ranges from 3 m to 2 m. The positioning service availability ranges from 95% to 99.9%. The scenarios of positioning service levels 1, 2, 3, and 5 require a 1s level latency, while positioning service levels 4 and 6 require a ms-level latency, which is only required in indoor scenarios of 5G enhanced positioning services, such as collaboration and collision avoidance of mobile robots and factory scenarios. For the relative positioning scenario of positioning service level 7, it is applicable to the positioning of two UEs within 10 m or the distance between a UE and a 5G positioning node within 10 m.
@@
When evaluating AI/ML based positioning, in addition to the common KPIs mentioned above, the performance of the AI model needs to be considered. For example, various overheads include open-air overhead, auxiliary information overhead, model delivery and transmission, etc. In addition, the complexity of the model’s reasoning needs to be considered, which includes the complexity of pre-processing and post-processing, the computational complexity in tera operations per second (TOPS), floating-point operations per second (FLOPS), and multiplication-accumulation operations (MACs), and the potential difference between actual complexity and evaluation complexity due to platform dependencies and optimization solutions. In addition, regardless of the underlying algorithm, model complexity should be reported in terms of the number of real-valued model parameters and operations. Finally, for model monitoring, performance indicators that need to be considered include the accuracy and relevance of monitoring indicators, related overhead, computational and memory complexity, and latency, which reflects the timeliness of monitoring results and response operations.
@@ IV-C1 Lifecycle Management for AI/ML models
  • Data Collection: The data collection function provides the input data required for model training, management, and inference. This data includes training data, monitoring data, and inference data, all essential for AI/ML model operation.
  • Model Training: The model training function executes the training, validation, and testing of AI/ML models. It may also generate performance metrics for model evaluation. Additionally, this function handles data preparation, such as preprocessing, cleaning, formatting, and transformation, using the training data provided by the data collection function.
  • Model Management: The model management function oversees the operation and monitoring of AI/ML models, including decisions related to model selection, activation, switching, and rollback. It ensures the correctness of inference operations based on data received from data collection and inference functions.
  • Model Inference: The inference function applies the AI/ML model to input data provided by the data collection function to produce outputs. This function also performs data preparation, such as preprocessing, when required, to ensure accurate inference results.
  • Model Storage: The model storage function stores trained or updated models, which can later be used for inference. This function also acts as a key point for model transfer/delivery processes and other protocol terminations.
These modules collectively address mechanisms for data collection, model training, updating, and sharing in wireless positioning, facilitating the effective deployment of AI/ML algorithms in wireless networks.
@@ IV-C2 AI/ML Models Deployment
In contrast, direct AI/ML positioning utilizes AI/ML models to determine the UE’s location directly. These models take raw wireless channel observations as inputs and estimate the UE’s position without relying on intermediate measurement steps. A prominent example of this approach is fingerprint-based positioning, where inputs such as channel impulse response (CIR) or power delay profile (PDP) are fed into an AI/ML model to directly estimate the UE’s location.
Based on the deployment location of AI/ML models within network entities and the distinction between direct and assisted positioning methods, 5 deployment cases have been identified. As shown in Fig. 9, the AI/ML models can be deployed on the UE, gNodeB (gNB), or the location management function (LMF), enabling flexibility in their application across different network architectures and positioning:
@@
    Case 1: In this case, the AI/ML model is deployed locally on the UE. The model can support both direct AI/ML positioning and AI/ML-assisted positioning using downlink PRS.
@@
    Case 2a: In this case, the AI/ML model is deployed on the UE. The UE utilizes the AI/ML model to perform measurements on the downlink PRS and transmits the measurement results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning.
@@
    Case 2b: In this case, the AI/ML model is deployed on the LMF. The UE performs measurements on the downlink PRS and transmits the results to the LMF, where the AI/ML model determines the UE’s location. This scenario only supports direct AI/ML positioning.
@@
    Case 3a: In this case, the AI/ML model is deployed on the gNB. The gNB uses the AI/ML model to perform measurements on the uplink SRS and transmits the results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning.
@@
    Case 3b: In this case, the AI/ML model is deployed on the LMF. The gNB performs measurements on the uplink SRS and transmits the results to the LMF, where the AI/ML model determines the UE’s position. This scenario only supports direct AI/ML positioning.
@@ IV-C3 Model Inputs and Outputs
For model training, training data can be generated by various network entities, including the UE, positioning reference unit (PRU), gNB, or LMF. For LMF-side model inference (Cases 2b and 3b), the input data is generated by the UE or gNB and terminates at the LMF. For gNB-side model inference (Case 3a), the input data is directly available within the gNB, reducing latency and ensuring efficient processing.
In cellular positioning, various measurements serve as critical inputs for AI/ML models to achieve accurate localization. These inputs are primarily derived from reference signals such as DL PRS and UL SRS, which are reused from existing 3GPP specifications but can be enhanced with AI-specific configurations. Key input types include high-dimensional information like delay profiles (DP), PDP, CIR, and CIR phase data. Additionally, timing-related measurements such as TOA, RSTD, RTT, and angle-related metrics extracted from CSI can also be leveraged as model inputs. Additionally, power metrics like DL PRS-RSRP and UL SRS-RSRP also contribute to improving positioning accuracy.

### v2

@@ IV Progress in AI Positioning Standards
In this section, we provide an overview of 3GPP positioning standards, from 1G to advanced AI/ML-supported techniques in 5G. We also discuss positioning KPIs, deployment scenarios, and future directions toward AI-native positioning in 6G.
@@ IV-A Evolution of 3GPP Standards for Positioning
1G networks used analog communication and lacked standardized positioning. Positioning was primarily based on signal strength (e.g., RSS) for BS selection and handovers. Some proprietary methods like TDOA and AOA in advanced mobile phone system (AMPS) signals, supported emergency and ITS applications [1, 188], but accuracy was limited.
2G introduced digital communication and standardized positioning in global system for mobile communications (GSM) [1], including Cell-ID, Time Advance, enhanced observed time difference (E-OTD), and assisted GPS (A-GPS) [189]. These improvements enabled services like E911 and laid the foundation for future enhancements.
3G technologies such as universal mobile telecommunications system (UMTS) and CDMA2000 introduced OTDOA, UTDOA, A-GPS [190], and radio frequency pattern matching (RFPM) [188]. CDMA2000 also used advanced forward link trilateration (AFLT) with A-GPS for better synchronization. These methods reduced positioning errors to 25–200 meters, addressing broader application needs.
In LTE Release 9, positioning became formally standardized. Key methods included [191]:
@@
To support cellular positioning, PRS signals were transmitted in dedicated subframes to minimize interference and enhance OTDOA performance. Furthermore, LTE positioning protocols, such as LTE positioning protocol (LPP), were defined to facilitate efficient communication between mobile terminals and positioning servers. Then, LTE-advanced (LTE-A) further enhanced positioning via UTDOA, improved OTDOA, RFPM [192], and advanced PRS performance. Additional techniques included D2D positioning, MIMO vertical estimation, and integration of Bluetooth, WLAN, and barometer data [193].
@@
With the evolution of 5G technologies, positioning capabilities in 3GPP standards have advanced significantly to meet the growing demands of diverse applications. In Release 16, the groundwork for 5G positioning was laid by defining positioning use cases and service requirements [194, 195]. These efforts were aimed at addressing the initial needs for positioning in 5G networks. Subsequent studies focused on enabling new radio (NR) positioning techniques across both FR1 (sub-6 GHz) and FR2 (mmWave) bands. Key aspects included the specification of NR Downlink and Uplink reference signals to support techniques such as Downlink TDOA (DL-TDOA), Downlink AOD (DL-AOD), Uplink TDOA (UL-TDOA), Uplink AOA (UL-AOA), Multi-cell RTT, and E-CID. Additionally, the NR positioning protocol A (NRPPa) was introduced [196], providing a standardized communication framework for positioning in 5G networks. To meet the demands of emerging 5G applications and vertical industries requiring higher accuracy, lower latency, and improved reliability, 3GPP launched the "Study on NR Positioning Enhancements" in Release 17 [197]. Key advancements included exploring positioning in in-coverage, partial coverage, and out-of-coverage scenarios to support diverse use cases, including sidelink-based positioning for V2X applications [198]. Release 18 introduced further improvements [199], such as PRS and sounding reference signal (SRS) bandwidth aggregation, NR carrier phase measurements, low-power high-accuracy support, and positioning for RedCap UEs. Sidelink positioning for V2X was also expanded.
@@
For AI/ML aspect, starting with Release 18, 3GPP initiated a study on AI/ML positioning [200]. This study explores the potential of AI/ML to improve positioning accuracy, details the general AI/ML framework, use cases for AI/ML positioning, and evaluation metrics and common KPIs for AI/ML positioning. These developments aim to complement traditional positioning methods by leveraging data-driven models to process large datasets and provide more accurate results.
TABLE V: 5G Positioning performance requirements
@@
Level 1 $10$ m 33 m N/A N/A $95$ % 1 s $\leq 30$ km/h (indoor) $\leq 250$ km/h (rural and urban) $\leq 500$ km/h (trains)
Level 2 33 m $99$ %
Level 3 $1$ m 22 m
@@
Level 4 11 m 22 m N/A N/A $99.9$ % 1515ms $\leq 30$ km/h indoor
Level 6 $0.3$ m 22 m 1010ms $\leq 30$ km/h indoor
@@ IV-B KPIs for Wireless Positioning
Positioning performance is assessed through a set of critical metrics that ensure the accuracy and reliability of location services. The 5G system is designed to deliver positioning services that meet the performance requirements outlined in Table. V. These requirements are inclusive of all UE types, including specialized UEs such as V2X and machine-type communication (MTC) devices.
@@
In 3GPP [195], 5G positioning use cases are divided into 5G positioning service area and 5G enhanced positioning service area. The former includes indoor and outdoor (rural and urban) scenarios. Indoor scenarios include locations inside buildings such as offices, hospitals, and industrial facilities. Outdoor scenarios support the positioning of vehicles and trains at speeds up to 250 km/h and 500 km/h. And the 5G enhanced positioning service area further supports dense urban areas (up to 60 km/h), vehicles in tunnels, and railway positioning. In addition to horizontal and vertical positioning accuracy, 3GPP Release 19 also introduces positioning service availability and positioning service latency to support mission-critical services. A total of 7 positioning service levels are defined, including 6 levels of absolute positioning accuracy requirements and 1 relative positioning accuracy requirement. In the absolute positioning accuracy requirements, the horizontal accuracy ranges from 10 m to 0.3 m, and the vertical accuracy ranges from 3 m to 2 m. The positioning service availability ranges from 95% to 99.9%. The scenarios of positioning service levels 1, 2, 3, and 5 require a 1s level latency, while positioning service levels 4 and 6 require a ms-level latency, which is only required in indoor scenarios of 5G enhanced positioning services, such as collaboration and collision avoidance of mobile robots and factory scenarios. For the relative positioning scenario of positioning service level 7, it is applicable to the positioning of two UEs within 10 m or the distance between a UE and a 5G positioning node within 10 m.
@@
When evaluating AI/ML-based positioning, in addition to the common KPIs mentioned above, the performance of the AI model needs to be considered. For example, various overheads include open-air overhead, auxiliary information overhead, model delivery and transmission, etc. The complexity of the model’s inference needs to be considered, which includes the complexity of pre-processing and post-processing, the computational complexity in tera operations per second (TOPS), floating-point operations per second (FLOPS), and multiplication-accumulation operations (MACs), and the potential difference between actual complexity and evaluation complexity due to platform dependencies and optimization solutions. In addition, regardless of the underlying algorithm, model complexity should be reported in terms of the number of real-valued model parameters and operations. Finally, for model monitoring, performance indicators that need to be considered include the accuracy and relevance of monitoring indicators, related overhead, computational and memory complexity, and latency, which reflects the timeliness of monitoring results and response operations.
@@ IV-C1 Lifecycle Management for AI/ML models
  • Data Collection: The data collection function provides the input data required for model training, management, and inference. This data includes training data, monitoring data, and inference data, all essential for AI/ML model operation. In the context of positioning, this module enables gNodeB (gNBs) to acquire rich and diverse data (e.g., CSI, RSRP) based on real-time network configuration and environmental conditions. These data are subsequently used for model training or inference. When more training data is needed, this module can also employ generative AI techniques such as GANs to augment datasets through synthetic data generation.
  • Model Training: The model training function executes the training, validation, and testing of AI/ML models. It may also generate performance metrics for model evaluation. Additionally, this function handles data preparation, such as preprocessing, cleaning, formatting, and transformation, using the training data provided by the data collection function. In positioning systems, this module can support techniques like transfer learning and meta learning to enable cross-domain model adaptation. Moreover, it is responsible for model compression and optimization, such as pruning or quantization, to reduce complexity while maintaining accuracy, enabling efficient deployment on resource-constrained devices.
  • Model Management: The model management function oversees the operation and monitoring of AI/ML models, including decisions related to model selection, activation, switching, and rollback. It ensures the correctness of inference operations based on data received from data collection and inference functions. This function plays a central role in managing multiple model versions customized for different environments and user KPI requirements. On one hand, it selects suitable models based on channel characteristics collected by the Data Collection module and the user-specific KPIs such as accuracy or latency. On the other hand, when degradation is detected (e.g., due to model aging), it initiates retraining or model replacement. Overall, this module coordinates interactions across the lifecycle pipeline to guarantee that positioning services meet real-time performance targets.
  • Model Inference: The inference function applies the AI/ML model to input data provided by the data collection function to produce outputs. This function also performs data preparation, such as preprocessing, when required, to ensure accurate inference results. This module executes the core inference process for positioning tasks, producing either location coordinates or intermediate outputs such as position-related signal parameters.
  • Model Storage: The model storage function stores trained or updated models, which can later be used for inference. This function also acts as a key point for model transfer/delivery processes and other protocol terminations. This component defines the structure of available models and manages their storage after training. It also tracks the complexity and accuracy of each stored model to support optimal model selection during runtime.
These modules collectively address mechanisms for data collection, model training, updating, and sharing in wireless positioning, facilitating the effective deployment of AI/ML algorithms in wireless networks. Specifically, the flexibility of the LCM framework provides the foundation for advanced AI capabilities such as transfer learning, online model adaptation, and network-aware configuration. These functions are expected to be vital for enabling intelligent and context-aware positioning services in 5G-Advanced and future 6G networks.
@@ IV-C2 AI/ML Models Deployment
TABLE VI: Comparison of AI/ML-Based Positioning Deployment Scenarios
Case 1 Case 2a Case 2b Case 3a Case 3b
Computation Location UE UE LMF gNB LMF
Communication Overhead Low Medium High Low Low
Signal Characteristics Downlink interference possible Downlink interference possible Downlink interference possible Lower transmit power than PRS Lower transmit power than PRS
Scalability Single-BS localization Single-/Multi-BS localization Single-/Multi-BS localization Single-BS localization Single-/Multi-BS localization
@@
In contrast, direct AI/ML positioning utilizes AI/ML models to determine the UE’s location directly. These models take raw wireless channel observations as inputs and estimate the UE’s position without relying on intermediate measurement steps. A prominent example of this approach is fingerprint-based positioning, where inputs such as CIR or power delay profile (PDP) are fed into an AI/ML model to directly estimate the UE’s location.
Based on the deployment location of AI/ML models within network entities and the distinction between direct and assisted positioning methods, 5 deployment cases have been identified. As shown in Fig. 7, the AI/ML models can be deployed on the UE, gNB, or the location management function (LMF), enabling flexibility in their application across different network architectures and positioning. In Fig. 7 and Table. VI, we provide a detailed description of each deployment scenario below.
@@
    Case 1: In this case, the AI/ML model is deployed locally on the UE. The model can support both direct AI/ML positioning and AI/ML-assisted positioning using downlink PRS. This deployment incurs low uplink communication overhead but requires significant computational resources on the UE side. It is less scalable for multi-BS scenarios and is thus better suited for single-BS localization. Moreover, the use of PRS may lead to downlink interference.
@@
    Case 2a: In this case, the AI/ML model is deployed on the UE. The UE utilizes the AI/ML model to perform measurements on the downlink PRS and transmits the measurement results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning. It offers moderate uplink communication overhead and good scalability for both single- and multi-BS localization. However, it still imposes high computational demands on the UE and may suffer from downlink PRS interference.
@@
    Case 2b: In this case, the AI/ML model is deployed on the LMF. The UE performs measurements on the downlink PRS and transmits the results to the LMF, where the AI/ML model determines the UE’s location. This setup supports direct AI/ML positioning and offers high scalability for both single- and multi-BS localization, while requiring small computational resources on the UE. The cost is the increased uplink communication overhead due to the raw signal transmission and the potential PRS interference in the downlink.
@@
    Case 3a: In this case, the AI/ML model is deployed on the gNB. The gNB uses the AI/ML model to perform measurements on the uplink SRS and transmits the results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning. It reduces the computational requirements on the UE side, but may result in limited participation from neighboring gNBs since SRS power is lower than PRS. In addition, since the AI model is deployed on the gNB, it is mainly applicable to single-BS scenarios.
@@
    Case 3b: In this case, the AI/ML model is deployed on the LMF. The gNB performs measurements on the uplink SRS and transmits the results to the LMF, where the AI/ML model determines the UE’s position. This scenario only supports direct AI/ML positioning. This approach minimizes the computational load on the UE and supports both single- and multi-BS positioning, providing good scalability. However, the relatively low power of the SRS may limit the number of participating gNBs, potentially affecting performance in sparse deployments.
@@ IV-C3 Model Inputs and Outputs
For model training, training data can be generated by various network entities, including the UE, positioning reference unit (PRU), gNB, or LMF. For LMF-side model inference (Cases 2b and 3b), the input data is generated by the UE or gNB and terminates at the LMF. For gNB-side model inference (Case 3a), the input data are directly available within the gNB, reducing latency and ensuring efficient processing.
In cellular positioning, various measurements serve as critical inputs for AI/ML models to achieve accurate localization. These inputs are primarily derived from reference signals such as DL PRS and UL SRS, which are reused from existing 3GPP specifications but can be enhanced with AI-specific configurations. Key input types include high-dimensional information like delay profiles (DP), PDP, CIR, and CIR phase data. Additionally, timing-related measurements such as TOA, RSTD, RTT, and angle-related metrics extracted from CSI can be leveraged as model inputs. Power metrics like DL PRS-RSRP and UL SRS-RSRP also contribute to improving positioning accuracy.
@@
IV-D Toward AI-Driven Positioning in 6G and Future Networks
Although the formal standardization process for 6G has not yet begun, it is widely anticipated that AI will play a central role in positioning services within AI-native network architectures. AI will no longer serve merely as an auxiliary tool, but will become a core component of next-generation network design. Based on this vision, we also discuss potential standardization directions for AI-based positioning technologies in future 6G networks in this section.
IV-D1 Native AI/ML Integration in Air Interface and Protocol Design
Unlike current 5G systems that rely on predefined mechanisms, such as PRS patterns, 6G is expected to adopt AI-native system designs, in which AI is deeply integrated not only into position estimation but also into broader aspects of air interface and protocol operation. Several standardization opportunities in this area include:
  • AI-Driven Pilot and Signaling Design: AI can be used to adaptively design reference signals and signaling strategies based on environmental conditions. For instance, dynamic adjustment of pilot density, beam configuration, or frequency band allocation can improve localization robustness in challenging environments such as THz communication [17], RIS-assisted networks [19], and ultra-dense deployments.
  • Joint Multi-Task AI Model Architectures: Future systems may adopt unified, multi-task AI models capable of jointly performing functions such as channel estimation, prediction, feedback compression, and positioning within a shared learning framework [169, 201]. This design can reduce model redundancy, improve computational efficiency, and enable task-aware inference pipelines.
  • AI-Based Protocol Adaptation: AI models can dynamically sense environmental variations (e.g., blockage, channel fluctuation, user mobility) and adapt key protocol parameters, such as scheduling periods, feedback granularity, and retransmission strategies, to optimize positioning accuracy and service continuity.
IV-D2 Multi-Source and Multi-Modal Positioning
6G systems are expected to natively support multi-source and multi-modal positioning by integrating signals from diverse infrastructures (such as cell-free networks [28], NTN [26], and RIS[19]) alongside heterogeneous sensory data. By combining radio signals from different access technologies with inputs from vision, LiDAR, inertial sensors, and map-based information, future systems could enable high-precision positioning and sensing applications. To realize these capabilities, the following aspects are expected to be addressed in future standardization efforts:
  • Unified Data Interfaces for Sensor and Network Fusion: Define standardized data formats and APIs for fusing radio and non-radio data sources across network elements and devices. This includes time synchronization, metadata structuring, and channel/measurement abstraction layers.
  • Scene-Aware Information Exchange Mechanisms: Establish standardized protocols for the exchange of environment-specific information (e.g., 3D maps and building layouts) among access points, edge servers, and user equipment. These mechanisms aim to enable collaborative localization and enhance situational awareness across heterogeneous networks. In addition, 6G can also consider the integration and synchronization of digital twin systems [202], which can provide dynamic, high-fidelity representations of physical environments to further improve positioning accuracy, consistency, and adaptability in complex scenarios
  • Cross-Domain Privacy-Preserving Learning Mechanisms: Develop standard-compliant methods to enable model training and update across different stakeholders (e.g., network operators, application vendors, and device manufacturers) without compromising data privacy or ownership.
IV-D3 AI-Specific Performance Metrics and QoS Requirements
Although existing 3GPP standards provide comprehensive KPIs for positioning accuracy, latency, these metrics are not tailored to AI-based localization algorithms and frameworks. As AI-driven positioning becomes a core enabler in 6G networks, it is essential to establish performance evaluation criteria that capture the unique characteristics, resource demands, and operational behavior of different AI models and deployment strategies. Future standardization should explicitly define the fundamental parameters of AI-driven positioning systems, including:
  • Network Resource Consumption: Quantify both communication overhead (e.g., channel and feature transmission) and computational resource usage (e.g., FLOPs, memory, inference time).
  • Data Transmission Volume: Measure the size of AI model inputs and outputs, including uplink/downlink data volume involved in distributed inference or collaborative learning [203].
  • Model Update Dynamics: Define metrics for training frequency, online adaptation cost, and update latency in dynamic or multi-user environments.
Building on these foundational parameters, AI-driven positioning systems should also be evaluated using a new class of AI-aware QoS indicators, including but not limited to:
  • Latency: It includes as end-to-end latency, transmission delay, and inference latency. Specifically, end-to-end latency refers to the total time from signal reception to position output, encompassing data preprocessing, feature extraction, model inference, and result post-processing.
  • Positioning Accuracy and Model Aging Rate: In addition to static positioning accuracy under standard conditions, it is important to evaluate the rate at which model performance degrades over time or across environments. This reflects the frequency at which models need to be retrained or updated to maintain acceptable accuracy.
  • Energy Efficiency: The energy consumption of the positioning process includes the computational energy in the inference process and the communication energy for data exchange.
IV-E Lessons Learned
The evolution of 3GPP standards reflects the growing maturity and significance of wireless positioning in cellular networks. Building upon traditional techniques, R18 marks a key milestone by introducing AI/ML frameworks into standardized positioning, enabling improved accuracy and robustness in complex scenarios such as NLOS conditions and dynamic environments. However, the integration of AI also brings new technical and standardization challenges, including the need for specialized performance metrics and comprehensive lifecycle management. The diversity of deployment options (across UE, gNB, and LMF), reveals inherent trade-offs between communication overhead, computational demands, and scalability. Furthermore, as positioning in 6G becomes tightly coupled with sensing and semantic understanding of the environment, standardization must also address challenges including AI-native network design, cross-modal data fusion, and privacy-preserving collaborative learning. This section highlights that the success of AI-driven positioning in future networks will rely not only on advances in algorithms but also on system-level orchestration and standardization frameworks tailored to the unique characteristics of AI models.